# Hölder curves with exotic tangent spaces

### v2

@@ Abstract.
An important implication of Rademacher’s Differentiation Theorem is that every Lipschitz curve $\Gamma$ infinitesimally looks like a line at almost all of its points in the sense that at $\mathcal{H}^{1}$-almost every point of $\Gamma$, the only tangent to $\Gamma$ is a straight line through the origin. In this article, we show that, in contrast, the infinitesimal structure of Hölder curves can be much more extreme. First we show that for every $s>1$ there exists a $(1/s)$-Hölder curve $\Gamma_{s}$ in a Euclidean space with $\mathcal{H}^{s}(\Gamma_{s})>0$ such that $\mathcal{H}^{s}$-almost every point of $\Gamma_{s}$ admits infinitely many topologically distinct tangents. Second, we study the tangents of self-similar connected sets (which are canonical examples of Hölder curves) and prove that the curves $\Gamma_{s}$ have the additional property that $\mathcal{H}^{s}$-almost every point of $\Gamma_{s}$ admits infinitely many homeomorphically distinct tangents to $\Gamma_{s}$ which are not admitted as (not even bi-Lipschitz to) tangents to any self-similar set at typical points.
@@ 1. Introduction
A major application of Rademacher’s Theorem is towards the understanding of the “infinitesimal structure” of Lipschitz curves (i.e. Lipschitz images of $[0,1]$). To state this application, let us first define the notion of tangents. Following [BL15], given a closed set $X\subset\mathbb{R}^{n}$ and a point $x\in X$, we say that a closed set $T$ is a tangent of $X$ at $x$ if there exists a sequence of positive scales $r_{j}$ that go to zero such that the blow-up sets $r_{j}^{-1}(X-x)$ converge to the set $T$ in the Attouch-Wets topology; see §2.4 for the precise definition. Other notions of metric space convergences which produce similar tangents are known in the literature; see [Gro81, Gro99, DS97]. Another well-known notion of infinitesimal structure in geometric measure theory is that of the tangent cone [Fed69, 3.1.21]. The tangent cone of $X$ at $x$ is the union of all tangents of $X$ at $x$, which means that some local information is lost. We denote by $\operatorname{Tan}(X,x)$ the collection of all tangents of $X$ at $x$. It is well-known that $\operatorname{Tan}(X,x)$ is nonempty and that if $T$ is in $\operatorname{Tan}(X,x)$, then $\lambda T$ is in $\operatorname{Tan}(X,x)$ for every $\lambda>0$.
@@
More generally, if $X$ is the Lipschitz (or even $p$-Sobolev for $p>n$) image of $0,1]^{n}$ into some $\mathbb{R}^{m}$ with $m>n$, then for $\mathcal{H}^{n}$-a.e. $x\in X$, the tangent space of $X$ at $x$ contains exactly one element which is an $n$-plane [BKV24]. Here and for the rest of this paper, given $\alpha>0$, we denote by $\mathcal{H}^{\alpha}(X)$ the Hausdorff $\alpha$-dimensional measure of a metric space $X$.
It is worthwhile to note that while typical points of a Lipschitz curve have a simple tangent space, exceptional points may exhibit extreme behaviors. In particular, there exists a Lipschitz curve $\Gamma\subset\mathbb{R}^{2}$ (or in any Euclidean space of dimension at least $2$) and a point $x_{0}\in\Gamma$ such that $\operatorname{Tan}(\Gamma,x_{0})$ contains every possible tangent; see Appendix A for a precise statement and the proof.
@@
Recent years have seen great interest in obtaining a Hölder version of the Analyst’s Traveling Salesperson Theorem, that is, in characterizing all sets which are contained in a Hölder curve; see for example [MM93, Rem98, MM00, BV17, BNV19, BZ20, BV21, BS23, SV24]. This notion of “Hölder rectifiability” is greatly motivated by the fact that many fractals in analysis on metric spaces admit a Hölder parameterization but not a Lipschitz one. For example, if the attractor of an iterated function system is connected (e.g. the von Koch curve, the Sierpiński gasket, the Sierpiński carpet), then it admits a Hölder parameterization by $[0,1]$; see [Rem98, BV21].
In lieu of the Lipschitz rectifiability results discussed above, a crucial step towards a comprehensive theory of Hölder rectifiable sets would be to understand the tangent spaces of Hölder curves at typical points. Unfortunately, unlike the Lipschitz case, it would be naive to expect that given $s>1$ there exists one single set $T_{s}$ which is the tangent of all $\frac{1}{s}$-Hölder curves at typical points. For example, consider the two self-similar sets $\mathcal{S}$ and $\mathcal{C}$ constructed as follows. Set $S_{0}=T$ where $T$ is an isosceles triangle with side-lengths $1,5^{-1/4},5^{-1/4}$, and for each $n\in\mathbb{N}$ set $S_{n}=\phi_{1}(S_{n-1})\cup\phi_{2}(S_{n-1})$ where $\phi_{1},\phi_{2}$ are the two similarities in the left image of Figure 1. The sets $S_{n}$ converge to a snowflake-like curve $\mathcal{S}=\bigcap_{n\in\mathbb{N}}S_{n}$. Similarly, set $C_{0}=[0,1]^{2}$ and for each $n\in\mathbb{N}$ set $C_{n}=\bigcup_{i=1}^{16}\psi_{i}(C_{n-1})$ where $\psi_{1},\dots,\psi_{16}$ are the similarities in the right image of Figure 1. The sets $C_{n}$ converge to a carpet-like set $\mathcal{C}=\bigcap_{n\in\mathbb{N}}S_{n}$. Both $\mathcal{S}$ and $\mathcal{C}$ are connected self-similar sets which have Hausdorff dimension equal to $s=\log{16}/\log{5}$ and positive $\mathcal{H}^{s}$ measure, and by a theorem of Remes [Rem98], both are $\frac{1}{s}$-Hölder curves. However, the tangents of $\mathcal{S}$ at $\mathcal{H}^{s}$-a.e. point are infinite snowflakes while the tangents of $\mathcal{C}$ at $\mathcal{H}^{s}$-a.e. point are infinite carpets; hence they are topologically different.
Figure 1. First stages of sets $\mathcal{S}$ (left) and $\mathcal{C}$ (right).
@@
For each $s>1$, there exists a $\frac{1}{s}$-Hölder curve $\Gamma_{s}$ in some Euclidean space such that $\mathcal{H}^{s}(\Gamma_{s})>0$ and for $\mathcal{H}^{s}$-almost every $x\in\Gamma_{s}$, the space $\operatorname{Tan}(\Gamma_{s},x)$ contains infinitely many topologically different tangents.
@@ 1.1. Tangents of self-similar sets
Perhaps the richest source of Hölder curves with positive and finite Hausdorff measure is that of self-similar sets. Recall that an iterated function system (abbv. IFS) of similarities on $\mathbb{R}^{n}$ is a finite collection $\mathcal{F}$ of contracting similarities of $\mathbb{R}^{n}$. We say that a set $K$ is the attractor of an IFS $\mathcal{F}$ if $K=\bigcup_{\phi\in\mathcal{F}}\phi(K)$. An IFS $\mathcal{F}$ satisfies the open set condition (abbv. OSC) if there exists a nonempty open set $U\subset\mathbb{R}^{n}$ such that for all distinct $\phi,\phi^{\prime}\in\mathcal{F}$ we have $\phi(U)\subset U$ and $\phi(U)\cap\phi^{\prime}(U)=\emptyset$. In this paper, we say that a set $Q$ is self-similar if it is the attractor of an IFS of similarities with the OSC. A well-known theorem of Hutchinson [Hut81] states that every self-similar set has finite and positive $\mathcal{H}^{s}$-measure for some $s>0$.
@@
In pursuit of the first of question, we investigate the tangents of self-similar sets and we show that if $Q$ is a self-similar set in a Euclidean space, then at $\mathcal{H}^{\operatorname{dim_{H}}(Q)}$-almost every point $x$ of $Q$, every tangent is locally made up of “big pieces of QQ”.
Proposition 1.3.
Let $\{\phi_{i}:\mathbb{R}^{N}\to\mathbb{R}^{N}\}_{i=1}^{m}$ be an IFS of similarities satisfying the OSC for some open set $U$ with $U\cap Q\neq\emptyset$, where $Q$ is the attractor, and let $s=\operatorname{dim_{H}}{Q}$. Then there exists $c_{0}\in\mathbb{N}$ depending only on the Lipschitz norms of $\phi_{1},\dots,\phi_{m}$ such that for $\mathcal{H}^{s}$-almost every $x\in Q$, every tangent $T\in\operatorname{Tan}(Q,x)$, and every $R>0$, there exist similarities $\{f_{i}^{R}:Q\to T\}_{i=1}^{M_{R}}$ with $M_{R}\leq c_{0}$ with Lipschitz norms in $c_{0}^{-1}R,R]$ such that
and for all distinct $i,j\in\{1,\dots,M_{R}\}$, $f_{i}^{R}(U)\cap f_{j}^{R}(U)=\emptyset$ and $f_{i}^{R}(U\cap Q)$ is an open subset of $T$.
Here and throughout we denote by $\bf 0$} the origin in $\mathbb{R}^{N}$. For the special case of self-similar sponges (including self-similar carpets such as the Sierpiński carpet), we show in Section 10 that at typical points of such a sponge, each tangent is locally self-similar, which means roughly that every open ball contains a similar copy of every other open ball as an open subset.
@@
Question 1.5 (Badger, Vellis).
@@
1.2. Outline of the proof of Theorem 1.2
@@
In Section 5 we show that, no matter what the choice function $\eta$ is, the resulting “limit set” $\Gamma_{s}$ will satisfy $0<\mathcal{H}^{s}(\Gamma_{s})<\infty$. Then, in Section 6, drawing inspiration from the techniques in parametrization results of attractors of iterated function systems (e.g. in [Hat85] and [BV21]), for each choice function $\eta$, we construct Lipschitz curves which approximate $\Gamma_{s}$ and converge to a $\frac{1}{s}$-Hölder parameterization of $\Gamma_{s}$.
@@ 2. Preliminaries
2.1. Dendrites
Lemma 2.1.
Let $T_{1},T_{2}$ be two dendrites in $\mathbb{R}^{n}$ that intersect on a point. Then $T_{1}\cup T_{2}$ is a dendrite.
Proof.
2.2. Words
For each even integer $n\geq 4$ we denote
$\mathcal{A}_{n}=\{1,\dots,5n-6\}.$
2.3. Combinatorial graphs and trees
$E\subset\{\{v,v^{\prime}\}:v,v^{\prime}\in V\text{ and }v\neq v^{\prime}\}.$
$\text{Im}(G):=\bigcup_{\{v,v^{\prime}\}\in E}[v,v^{\prime}]$
A simple path in $G$ is a set $\gamma=\{\{v_{1},v_{2}\},\dots,\{v_{n-1},v_{n}\}\}\subset E$ such that for all distinct $i,j\in\{1,\dots,n\}$ we have $v_{i}\neq v_{j}$; in this case we say that $\gamma$ joins $v_{1}$, $v_{n}$. A graph $T=(V,E)$ is a combinatorial tree if for any distinct $v,v^{\prime}$ there exists a unique simple path $\gamma$ that joins $v$ with $v^{\prime}$.
2.4. Tangents
If $A$ and $B$ are nonempty subsets of a Euclidean space $\mathbb{R}^{N}$, then we define the excess of A over B as
@@ 2.4. Tangents
see [Bee93, §3.1], [BL15], and [BET17, Appendix A]. In the next remark, we list six important properties of the excess which will be used heavily throughout this article. The first and fifth properties follow straightforwardly from the definition of excess while the other four are given in [BL15, Section 2].
@@
Lemma 2.3 ([Bee93, Chapter 3]).
There exists a metrizable topology, called the Attouch-Wets topology, on $\mathfrak{C}(\mathbb{R}^{N})$ in which a sequence of sets $(X_{i})_{i=1}^{\infty}\subset\mathfrak{C}(\mathbb{R}^{N})$ converges to a set $X\in\mathfrak{C}(\mathbb{R}^{N})$ if and only if for every $r>0$,
@@
Moreover, the subcollection $\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ is sequentially compact; that is, for any sequence $(X_{i})_{i=1}^{\infty}\subset\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ there exists a subsequence $(X_{i_{j}})_{j=1}^{\infty}$ and a set $X\in\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ such that $(X_{i_{j}})_{j=1}^{\infty}$ converges to $X$.
@@
If $(A_{j})_{j\in\mathbb{N}}\subset\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ is a sequence converging to a set $A\in\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ with respect to the Attouch-Wets topology, then for every $r>0$ and each $\epsilon>0$, there exists an integer $j_{0}\in\mathbb{N}$ so that for every pair of integers $j_{1},j_{2}\geq j_{0}$, $\operatorname{excess}(A_{j_{1}}\cap\overline{B}({\bf 0},r),A_{j_{2}})<\epsilon$.
@@
Let $\epsilon>0$ and let $r>0$. Since $A_{j}\to A$ in the Attouch-Wets topology as $j\to\infty$, there exists an integer $j_{0}\in\mathbb{N}$ so that for every integer $j\geq j_{0}$
@@
Now since $j_{1}\geq j_{0}$, $\operatorname{excess}(A_{j_{1}}\cap\overline{B}({\bf 0},r),A)<\min\{\epsilon/2,r\}$, so for every $x\in A_{j_{1}}\cap\overline{B}({\bf 0},r)$ there exists a point $y\in A$ so that $|x-y|<\min\{\epsilon/2,r\}$. Then it must hold that such a point $y$ is contained in $\overline{B}({\bf 0},2r)\cap A$, and therefore $\operatorname{excess}(A_{j_{1}}\cap\overline{B}({\bf 0},r),A\cap\overline{B}({\bf 0},2r))<\epsilon/2$. Additionally, since $j_{2}\geq j_{0}$, we have also that $\operatorname{excess}(A\cap\overline{B}({\bf 0},2r),A_{j_{2}})<\epsilon/2$, and this completes the proof. ∎
Let $X\in\mathfrak{C}(\mathbb{R}^{N})$ and let $x\in X$. Following [BL15], we say that a set $T\in\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ is a tangent set of $X$ at $x$ if there exists a sequence of scales $(r_{m})_{m\in\mathbb{N}}>0$ such that $r_{m}\to 0$ and $\frac{X-x}{r_{m}}\to T$ with respect to the Attouch-Wets topology. We denote by $\operatorname{Tan}(X,x)$ the set of all tangent sets to $X$ at $x$.
For the next lemma, denote by $\mathfrak{C}_{U}(\mathbb{R}^{N};{\bf 0})$ the collection of all sets $X\in\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ such that every component of $X$ is unbounded.
@@
If $X\subset\mathbb{R}^{N}$ is a nondegenerate continuum and $x\in X$, then $\operatorname{Tan}(X,x)\subset\mathfrak{C}_{U}(\mathbb{R}^{N};{\bf 0})$.
@@
Fix $T\in\operatorname{Tan}(X,x)$. First, we prove that the component of $T$ which contains $\bf 0$} is unbounded. To this end, let $(r_{j})>0$ be a sequence of scales converging to 0 so that $(r_{j})^{-1}(X-x)\to T$ as $j\to\infty$ in the Attouch-Wets topology. Given $j\in\mathbb{N}$, define the similarity $g_{j}:X\to\mathbb{R}^{N}$ by $g_{j}(y)=(r_{j})^{-1}(y-x)$, and for $R>0$ let $Q_{j,R}$ be the component of $g_{j}(X)\cap\overline{B}({\bf 0},R)$ containing $\bf 0$}. Since $X$ is connected and nondegenerate, for each $R>0$ there exists a $j_{R}\in\mathbb{N}$ so that $Q_{j,R}\cap\partial B({\bf 0},R)\neq\emptyset$ for every integer $j\geq j_{R}$. Note that for every pair of scales $R_{2}>R_{1}>0$ and for each $j\in\mathbb{N}$, $Q_{j,R_{1}}\subset Q_{j,R_{2}}$. By compactness of $\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$, for each $R>0$ the sequence $(Q_{j,R})_{j\in\mathbb{N}}$ has an Attouch-Wets subsequential limit $K_{R}$, and for $R_{2}>R_{1}>0$ we have $K_{R_{1}}\subset K_{R_{2}}$. Additionally, since $g_{j}(X)$ converges to $T$ in the Attouch-Wets topology, we have that $K_{R}\subset T$, ${\bf 0}\in K_{R}$, and $K_{R}\cap\partial B({\bf 0},R)\neq\emptyset$ for every $R>0$. Therefore, $\bigcup_{m\in\mathbb{N}}K_{m}\subset T$ and the former union is unbounded, connected, and contains $\bf 0$}, so the component of $T$ containing $\bf 0$} is unbounded, thus $T$ has an unbounded component.
Assume now for a contradiction that $T$ has a bounded component $K$. Let $L$ be the component of $T^{\prime}=T\cup\{\infty\}$ in $\mathbb{S}^{N}$ that contains $\infty$. Then clearly $K,L$ are distinct connected components of $T^{\prime}$. Regarding $K$ and $L$ as quasi-components of $T^{\prime}$, one can choose two disjoint closed sets $Y,Z\subset T^{\prime}$ satisfying $K\subset Y$, $L\subset Z$, and $T^{\prime}=Y\cup Z$. Then both $Y$ and $Z$ are closed in $\mathbb{S}^{N}$, and thus by normality there exists an open set $U$ such that
@@
Fix $R>0$ large enough that $Y\subset\overline{B}({\bf 0},R)$ and $Z\cap\overline{B}({\bf 0},R)\neq\emptyset$. By definition of Attouch-Wets convergence, there exists a $j_{0}\in\mathbb{N}$ so that for every integer $j\geq j_{0}$,

### v3

@@ Abstract.
An important implication of Rademacher’s Differentiation Theorem is that every Lipschitz curve $\Gamma$ infinitesimally looks like a line at almost all of its points in the sense that at $\mathcal{H}^{1}$-almost every point of $\Gamma$, the only tangent to $\Gamma$ is a straight line through the origin. In this article, we show that, in contrast, the infinitesimal structure of Hölder curves can be much more extreme. First, we show that for every $s>1$ there exists a $(1/s)$-Hölder curve $\Gamma_{s}$ in a Euclidean space with $\mathcal{H}^{s}(\Gamma_{s})>0$ such that $\mathcal{H}^{s}$-almost every point of $\Gamma_{s}$ admits infinitely many topologically distinct tangents. Second, we study the tangents of self-similar connected sets (which are canonical examples of Hölder curves) and prove that the curves $\Gamma_{s}$ have the additional property that $\mathcal{H}^{s}$-almost every point of $\Gamma_{s}$ admits infinitely many homeomorphically distinct tangents to $\Gamma_{s}$ which are not admitted as (not even bi-Lipschitz to) tangents to any self-similar set at typical points.
@@ 1. Introduction
A major application of Rademacher’s Theorem is towards the understanding of the “infinitesimal structure” of Lipschitz curves (i.e. Lipschitz images of $[0,1]$). To state this application, let us first define the notion of tangents. Following [BL15], given a closed set $X\subset\mathbb{R}^{n}$ and a point $x\in X$, we say that a closed set $T$ is a tangent of $X$ at $x$ if there exists a sequence of positive scales $r_{j}$ that go to zero such that the blow-up sets $r_{j}^{-1}(X-x)$ converge to the set $T$ in the Attouch-Wets topology; see §2.1 for the precise definition. Other notions of metric space convergences which produce similar tangents are known in the literature; see [Gro81, Gro99, DS97, Fur14]. Another well-known notion of infinitesimal structure in geometric measure theory is that of the tangent cone [Fed69, 3.1.21]. The tangent cone of $X$ at $x$ is the union of all tangents of $X$ at $x$, which means that some local information is lost. We denote by $\operatorname{Tan}(X,x)$ the collection of all tangents of $X$ at $x$. It is well-known that $\operatorname{Tan}(X,x)$ is nonempty and that if $T$ is in $\operatorname{Tan}(X,x)$, then $\lambda T$ is in $\operatorname{Tan}(X,x)$ for every $\lambda>0$.
@@
Here and for the rest of this paper, given $\alpha>0$, we denote by $\mathcal{H}^{\alpha}(X)$ the Hausdorff $\alpha$-dimensional measure of a metric space $X$. It is worthwhile to note that while typical points of a Lipschitz curve have a simple tangent space, exceptional points may exhibit extreme behaviors. In particular, there exists a Lipschitz curve $\Gamma\subset\mathbb{R}^{2}$ (or in any $\mathbb{R}^{N}$ with $N\geq 2$) and a point $x_{0}\in\Gamma$ such that $\operatorname{Tan}(\Gamma,x_{0})$ contains every possible tangent; see Appendix A for a precise statement and the proof.
@@
Recent years have seen great interest in obtaining a Hölder version of the Analyst’s Traveling Salesperson Theorem, that is, in characterizing all sets which are contained in a Hölder curve; see for example [MM93, Rem98, MM00, BV19, BNV19, BZ20, BV21, SV24, BS25]. This notion of “Hölder rectifiability” is greatly motivated by the fact that many fractals in analysis on metric spaces admit a Hölder parameterization but not a Lipschitz one. For example, if the attractor of an iterated function system is connected (e.g. the von Koch curve, the Sierpiński gasket, the Sierpiński carpet), then it admits a Hölder parameterization by $[0,1]$; see [Rem98, BV21].
In lieu of the Lipschitz rectifiability results discussed above, a crucial step towards a comprehensive theory of Hölder rectifiable sets would be to understand the tangent spaces of Hölder curves at typical points. Unfortunately, unlike the Lipschitz case, it would be naive to expect that given $s>1$ there exists one single set $T_{s}$ which is the tangent of all $\frac{1}{s}$-Hölder curves at typical points. For example, let $\mathcal{S}$ be a self-similar Koch-like snowflake of dimension $s=\log{16}/\log{5}$ and let $\mathcal{C}$ be a Sierpiński-type carpet obtained by dividing the unit square into 25 pieces and removing the 9 center squares; see Figure 1. Both $\mathcal{S}$ and $\mathcal{C}$ are connected self-similar sets which have Hausdorff dimension equal to $s=\log{16}/\log{5}$ and positive $\mathcal{H}^{s}$ measure, and by a theorem of Remes [Rem98], both are $\frac{1}{s}$-Hölder curves. However, the tangents of $\mathcal{S}$ at $\mathcal{H}^{s}$-a.e. point are infinite snowflakes while the tangents of $\mathcal{C}$ at $\mathcal{H}^{s}$-a.e. point are infinite carpets; hence they are topologically different.
Figure 1. Snowflake $\mathcal{S}$ (left) and carpet $\mathcal{C}$ (right).
@@
For each $s>1$, there exists a $\frac{1}{s}$-Hölder curve $\Gamma_{s}$ in some Euclidean space such that $\mathcal{H}^{s}(\Gamma_{s})>0$ and for $\mathcal{H}^{s}$-almost every $x\in\Gamma_{s}$, the space $\operatorname{Tan}(\Gamma_{s},x)$ contains infinitely many tangents which are pairwise not homeomorphic.
@@ 1.1. Tangents of self-similar sets
Perhaps the richest source of Hölder curves with positive and finite Hausdorff measure is that of self-similar sets. In this paper, we say that a nonempty compact set $Q\subset\mathbb{R}^{N}$ is self-similar if there exist contracting similarities $\phi_{1},\dots,\phi_{k}$ on $\mathbb{R}^{N}$ and there exists a nonempty open set $U\subset\mathbb{R}^{N}$ such that $Q=\phi_{1}(Q)\cup\cdots\cup\phi_{k}(Q)$ and for all distinct $i,j$ we have $\phi_{i}(U)\subset U$ and $\phi_{i}(U)\cap\phi_{j}(U)=\emptyset$. A well-known theorem of Hutchinson [Hut81] states that every self-similar set has finite and positive $\mathcal{H}^{s}$-measure for some $s>0$.
@@
Towards the first question, we show that if $Q\subset\mathbb{R}^{N}$ is a self-similar set, then at $\mathcal{H}^{\operatorname{dim_{H}}(Q)}$-almost every point $x$ of $Q$, every tangent $T$ is locally made up of large copies of $Q$: there exists a controlled number of rescaled copies $Q_{1},\dots,Q_{n}$ of $Q$, of diameters comparable to $R$, that are “essentially disjoint” and
$\overline{B}({\bf 0},R)\cap T\subset Q_{1}\cup\cdots\cup Q_{n}\subset\overline{B}({\bf 0},2R)\cap T$
where $\bf 0$} denotes the origin in $\mathbb{R}^{N}$. See Proposition 10.1 for the precise statement. Furthermore, for the special case of self-similar sponges (including self-similar carpets such as the Sierpiński carpet), we show in Section 10 that at typical points of such a sponge, each tangent is locally self-similar, which means roughly that every open ball contains a similar copy of every other open ball as an open subset.
@@
Question 1.4 (Badger, Vellis).
@@
1.2. Outline of the proofs of main results
@@
In Section 5 we show that, no matter what the choice function $\eta$ is, the resulting “limit set” $\Gamma_{s}$ will satisfy $0<\mathcal{H}^{s}(\Gamma_{s})<\infty$.
In Section 6, perhaps the most technical part of this paper, drawing inspiration from the techniques in parameterization results of attractors of iterated function systems [BV21], we construct, for each choice function $\eta$, a sequence of Lipschitz curves which approximate $\Gamma_{s}$ and converge to a $\frac{1}{s}$-Hölder parameterization of $\Gamma_{s}$. We note that Remes’s theorem [Rem98] asserts the existence of a Hölder parameterization for self-similar sets but does not give the parameterization itself. Constructions of such parameterizations exist in the literature [RZ16, DRZ19] for special cases of self-similar sets (including self-similar carpets). Our approach in Section 6 can be used to construct such parameterizations for statistically self-similar carpets and sponges.
@@
Acknowledgements
We thank the anonymous referees for their valuable comments which greatly improved the exposition of the paper.
@@ 2. Preliminaries
Given two points $a,b\in\mathbb{R}^{N}$, we denote by |ab||a-b| the Euclidean distance between these points. Given a point $a\in\mathbb{R}^{N}$ and a set $X\subset\mathbb{R}^{N}$ we write $\operatorname{dist}(a,X):=\inf_{x\in X}|a-x|$. Finally, given sets $X,Y\subset\mathbb{R}^{N}$, we write
$\operatorname{dist}(X,Y):=\inf_{x\in X,y\in Y}|x-y|=\inf_{x\in X}\operatorname{dist}(x,Y)=\inf_{y\in Y}\operatorname{dist}(y,X).$
Furthermore, given nonempty sets $A,B\subset\mathbb{R}^{N}$, we define the excess of A over B as
@@
see [Bee93, §3.1], [BL15], and [BET17, Appendix A]. We define also the Hausdorff distance between $A$ and $B$ by
$\operatorname{dist_{H}}(A,B):=\max\left\{\operatorname{excess}(A,B),\operatorname{excess}(B,A)\right\}.$
Unlike the Hausdorff distance, excess is not symmetric. Clearly, $\operatorname{excess}(A,B)\leq\operatorname{dist_{H}}(A,B)$.
In the next remark, we list six important properties of the excess which will be used heavily throughout this article. The first and fifth properties follow straight from the definition while the other four are given in [BL15, Section 2].
@@ 2.1. Tangents
Lemma 2.2 ([Bee93, Theorem 3.1.7], [DS97, Lemma 8.2]).
There exists a metrizable topology on $\mathfrak{C}(\mathbb{R}^{N})$, called the Attouch-Wets topology, in which a sequence of sets $(X_{m})_{m=1}^{\infty}\subset\mathfrak{C}(\mathbb{R}^{N})$ converges to a set $X\in\mathfrak{C}(\mathbb{R}^{N})$ if and only if for every $r>0$,
@@
Moreover, the subcollection $\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ is sequentially compact; that is, for any sequence $(X_{m})_{m=1}^{\infty}\subset\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ there exists a subsequence $(X_{m_{j}})_{j=1}^{\infty}$ and a set $X\in\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ such that $(X_{m_{j}})_{j=1}^{\infty}$ converges to $X$.
@@
If $(A_{j})_{j\in\mathbb{N}}\subset\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ is a sequence converging to a set $A\in\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ with respect to the Attouch-Wets topology, then for every $r>0$ and each $\epsilon>0$, there exists an integer $j_{0}\in\mathbb{N}$ such that for every pair of integers $j_{1},j_{2}\geq j_{0}$, $\operatorname{excess}(A_{j_{1}}\cap\overline{B}({\bf 0},r),A_{j_{2}})<\epsilon$.
@@
Let $\epsilon>0$ and let $r>0$. Since $A_{j}\to A$ in the Attouch-Wets topology as $j\to\infty$, there exists an integer $j_{0}\in\mathbb{N}$ such that for every integer $j\geq j_{0}$
@@
Now since $j_{1}\geq j_{0}$, $\operatorname{excess}(A_{j_{1}}\cap\overline{B}({\bf 0},r),A)<\min\{\epsilon/2,r\}$, thus for every $x\in A_{j_{1}}\cap\overline{B}({\bf 0},r)$ there exists a point $y\in A$ such that $|x-y|<\min\{\epsilon/2,r\}$. Then it must hold that such a point $y$ is contained in $\overline{B}({\bf 0},2r)\cap A$, and therefore $\operatorname{excess}(A_{j_{1}}\cap\overline{B}({\bf 0},r),A\cap\overline{B}({\bf 0},2r))<\epsilon/2$. Additionally, since $j_{2}\geq j_{0}$, we have also that $\operatorname{excess}(A\cap\overline{B}({\bf 0},2r),A_{j_{2}})<\epsilon/2$, which completes the proof. ∎
Definition 2.4 ([BL15, Definition 3.1]).
Let $X\in\mathfrak{C}(\mathbb{R}^{N})$ and let $x\in X$. We say that a set $T\in\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ is a tangent set of $X$ at $x$ if there exists a sequence of scales $(r_{m})_{m\in\mathbb{N}}>0$ such that $r_{m}\to 0$ and $\frac{X-x}{r_{m}}\to T$ with respect to the Attouch-Wets topology. We denote by $\operatorname{Tan}(X,x)$ the set of all tangent sets of $X$ at $x$.
Since $\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ is sequentially compact, if $X\in\mathfrak{C}(\mathbb{R}^{N})$ and if $x_{0}\in X$, then the sequence of sets $(n(X-x_{0}))_{n\in\mathbb{N}}$ has a subsequential limit in $\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$ and such a limit must be a tangent of $X$ at $x_{0}$. Therefore, for any $X\in\mathfrak{C}(\mathbb{R}^{N})$ and any $x_{0}\in X$, we have that $\emptyset\neq\operatorname{Tan}(X,x_{0})\subset\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$.
For a notion of tangents with respect to Hausdorff metric (microsets) see [Fur14] and see [Buc03, BR06] for microsets of continuous functions.
@@
If $X\subset\mathbb{R}^{N}$ is a nondegenerate continuum, $x\in X$, and $T\in\operatorname{Tan}(X,x)$, then each component of $T$ is unbounded.
@@
Fix $T\in\operatorname{Tan}(X,x)$. First, we prove that the component of $T$ which contains $\bf 0$} is unbounded. To this end, let $(r_{j})>0$ be a sequence of scales converging to 0 such that $(r_{j})^{-1}(X-x)\to T$ as $j\to\infty$ in the Attouch-Wets topology. Given $j\in\mathbb{N}$, define the similarity $g_{j}:X\to\mathbb{R}^{N}$ by $g_{j}(y)=(r_{j})^{-1}(y-x)$, and for $R>0$ let $Q_{j,R}$ be the component of $g_{j}(X)\cap\overline{B}({\bf 0},R)$ containing $\bf 0$}. Since $X$ is connected and nondegenerate, for each $R>0$ there exists a $j_{R}\in\mathbb{N}$ such that $Q_{j,R}\cap\partial B({\bf 0},R)\neq\emptyset$ for every integer $j\geq j_{R}$. Note that for every pair of scales $R_{2}>R_{1}>0$ and for each $j\in\mathbb{N}$, $Q_{j,R_{1}}\subset Q_{j,R_{2}}$. By compactness of $\mathfrak{C}(\mathbb{R}^{N};{\bf 0})$, for each $R>0$ the sequence $(Q_{j,R})_{j\in\mathbb{N}}$ has an Attouch-Wets subsequential limit $K_{R}$, and for $R_{2}>R_{1}>0$ we have $K_{R_{1}}\subset K_{R_{2}}$. Additionally, since $g_{j}(X)$ converges to $T$ in the Attouch-Wets topology, we have that $K_{R}\subset T$, ${\bf 0}\in K_{R}$, and $K_{R}\cap\partial B({\bf 0},R)\neq\emptyset$ for every $R>0$. Therefore, $\bigcup_{m\in\mathbb{N}}K_{m}\subset T$ and the former union is unbounded, connected, and contains $\bf 0$}, hence the component of $T$ containing $\bf 0$} is unbounded, and thus $T$ has an unbounded component.
Assume now for a contradiction that $T$ has a bounded component $K$ and let $\mathbb{S}^{N}$ denote the 1–point compactification of $\mathbb{R}^{N}$. Let $L$ be the component of $T^{\prime}=T\cup\{\infty\}$ in $\mathbb{S}^{N}$ that contains $\infty$. Then clearly $K,L$ are distinct connected components of $T^{\prime}$. Regarding $K$ and $L$ as quasi-components of $T^{\prime}$, one can choose two disjoint closed sets $Y,Z\subset T^{\prime}$ satisfying $K\subset Y$, $L\subset Z$, and $T^{\prime}=Y\cup Z$. Then both $Y$ and $Z$ are closed in $\mathbb{S}^{N}$, and thus by normality there exists an open set $U$ such that
@@
Fix $R>0$ large enough that $Y\subset\overline{B}({\bf 0},R)$ and $Z\cap\overline{B}({\bf 0},R)\neq\emptyset$. By definition of Attouch-Wets convergence, there exists a $j_{0}\in\mathbb{N}$ such that for every integer $j\geq j_{0}$,